The sample contains 9 E15 electrons in volume 6 (mm^3). What is the electron concentration in the sample? (Unit: 1/(m^3))
Answer #1

No of electron in sample=9E15=9×10^-15 and sample volume=6mm^3=6×10^-3m

Electron Conentration=(No. Of Electron)/(Volume) = (9×10^-15)/(6×10^-3) =1.5×10^-12 /m^3